ive tried this plenty of times i get stuck because i know how to solve the problom when there are two vairibles but not three please helpp! i have a test on this tommorow
the question on my worksheet is write a quadratic function in standard form whose graph passes through the given points:
(1,-2),(-2,1),(3,6)

Answer by stanbon(75887)

write a quadratic function in standard form whose graph passes through the given points:
(1,-2),(-2,1),(3,6)
------------------------------
Form: ax^2 + bx + c = y
Substitute each point for the x and the y values
to get 3 equations. Then solve the system for a,b,c.
--------------------------------
Using(1,-2) you get: a + b + c = -2
Using(-2,1) you get:4a -2b + c = 1
Using(3,6) you get:9x^2+3b + c = 6
----------------------------------------
Solve by any method you know to get:
a = 1
b = 0
c = -3
===============
Cheers,
Stan H.